OpenStudy (anonymous):

The sum of two consecutive integers is at least 36. What is the least possible pair of integers? A. 18 and 19 B. 19 and 20 C. 17 and 18 D. 16 and 17

OpenStudy (anonymous):

@KirbyLegs when I see your little picture pop up I'm like "Oh, It's my homie!!! xDDDD "

OpenStudy (anonymous):

Add these numbers up right? A: 18+19 = 37 B: 19+20 = 39 C: 17+18 = 35 D:16 + 17 = 33 The answer is A p.s. hey itsanna!

OpenStudy (anonymous):

Yay! Okay that's easy... :D

OpenStudy (anonymous):

16+17 = 33 17+18 = 35 18+19 = 37 19+20 = 39 So The first two ones don't add up to 36, so they are out. Next you are left with 18+19=37 and 19+20=39 So 37<39 That means 18+19 < 19+20 So the smallest pair of integers that add up to at least 36 is 18+19.